OpenStudy (anonymous):

least common multiple 20 6

OpenStudy (anonymous):

60

OpenStudy (anonymous):

60 ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

cant be 60

OpenStudy (turingtest):

yes 60

OpenStudy (anonymous):

so i dont get the question

OpenStudy (anonymous):

Again look at the factors of each: 6: 2, 3 20: 2, 2, 5 Multiply them all 2*3*2*5=60

OpenStudy (across):

20=(2)(10)=(2)(2)(5) 6=(2)(3) (2)(2)(3)(5)=60

OpenStudy (turingtest):

3*20=6*10=60

OpenStudy (anonymous):

If it's a question on least common multiple then it is in fact 60

hero (hero):